If you look at stock prices over any year, you will find a high and low stock price for the year. Instead of a single benchmark PE ratio, we now have a high and low PE ratio for each year. We can use these ratios to calculate a high and a low stock price for the next year. Suppose we have the following information on a particular company over the past four years:

Year 1 Year 2 Year 3 Year 4
  High price $ 98.60 $ 122.20 $ 131.60 $ 148.23
  Low price 73.43 89.54 70.22 116.75
  EPS 7.88 9.63 10.71 12.10

  

Earnings are projected to grow at 4 percent over the next year.
a.

What is your high target stock price over the next year? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

b. What is your low target stock price over the next year? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Computation of earnings per share for next year.

Current year EPS = $ 12.10

Growth rate = 4%

Next year EPS = Current year EPS x (1+growth rate)

Next year EPS = 12.10(1.04) = $ 12.584

Computation of the high PE ratio and low PE ratio for each year.

PE ratio = Stock price / EPS

Year High PE Ratio Low PE Ratio
1 98.60/7.88 12.512 73.43/7.88 9.318
2 122.20/9.63 12.689 89.54/9.63 9.298
3 131.60/10.71 12.287 70.22/10.71 6.556
4 148.23/12.10 12.250 116.75/12.10 9.648
Total 49.738 34.82

Computation of average high PE ratio and average low PE ratio.

Average high PE ratio = 49.738/ 4 = 12.434

Average low PE ratio = 34.82/4 = 8.705

(a) Computation of high target stock price over the next year.

High target stock price = Average high PE ratio x Expected EPS for next year

High target stock price = 12.434 x 12.584 = $ 156.46

(b) Computation of low target stock price over the next year.We have,

Low target stock price = Average low PE ratio x Expected EPS for next year

Low target stock price = 8.705 x 12.584 = $ 109.54